PTUURBIJ + 5. Velocity field of a 2-dimensional flow motion of an inviscid and incompressible fluid is given by, u=x, v=y,
Answer #1

Write the velocity equation in vector form. ū = uî +vſ+ wk=(x)i +(y)+(0)ế=xi+y?ſ +O Substitute the given coordinate values. V=C2)(*]+(3) L]+020] =(x®).0+(yº)-3y2 +(0) 0 = 0 +3y3 +0 =a, = 3y3 Calculate the z-component of acceleration. =(29)[0+(30) ,0